OpenStudy (anonymous):

4/3+2/x-5=-6/3x-15 x = –8 x = –2 x = 2 x = 8

OpenStudy (anonymous):

\[\frac{4}{3}+\frac{2}{x}-5 = -\frac{6}{3x}-15\]?

OpenStudy (anonymous):

(4/3)+(2/x-5)=-6/(3x-15)

OpenStudy (anonymous):

4/3+2/(x-5)=-6/(3x-15) there we go lol

OpenStudy (anonymous):

multiply through by 3x 4x + 6 - 5 = -6 - 45x 49x = -7 x = -1/7

OpenStudy (anonymous):

oops soorry = error in first line 4x + 6 - 15x = -6 - 45x 34x = -12 x = -12/34 = -6/17